OpenStudy (anonymous):

I give medals to the person that explains how to do it without just giving me the Answer! Eliza started her savings account with $100. Each month she deposits $25 into her account. Write the function that describes the relationship between the amount Eliza deposits and the total amount in her account. Then, determine the average rate of change in the account from the 2nd month to the 10th month

OpenStudy (lexi724):

i messaged you

OpenStudy (lexi724):

Average rate of change formula: m=y2−y1x2−x1 y2=350 y1=150 x2=10 x1=2 m=350−15010−2 m=2008 m=25

OpenStudy (lexi724):

uga formulas wont show

OpenStudy (anonymous):

hi

OpenStudy (anonymous):